### v1

# Dynamic Level Sets

Abstract
A mathematical concept is identified and analyzed that is implicit in the 2012 paper Turing Incomputable Computation [6], presented at the Alan Turing Centenary Conference (Turing 100, Manchester). The concept, called dynamic level sets, is distinct from mathematical concepts in the standard literature on dynamical systems, topology, and computability theory. A new mathematical object is explained and why it may have escaped prior characterizations, including the classical result of de Leeuw, Moore, Shannon, and Shapiro [4] that probabilistic Turing machines compute no more than deterministic ones.
1 Level Sets in Pure Mathematics
In standard analysis, topology, and dynamical systems theory, a level set of a real-valued function $f:X\to\mathbb{R}$ at value $c$ is the preimage
$f^{-1}(c)\;=\;\{\,x\in X\mid f(x)=c\,\}.$
Closely related are the sublevel sets $f^{-1}((-\infty,c])=\{x\mid f(x)\leq c\}$.
While contour lines [^1] Charles Hutton used them in his 1774 survey of Schiehallion mountain in Scotland. on maps [10] (curves of equal elevation) represent the earliest graphical use of the concept, the first systematic mathematical use of level sets in dynamics appears in Lyapunov’s 1892 doctoral dissertation [12]. Lyapunov’s stability analysis depends on the geometry of sublevel sets of a positive-definite function $V$: the invariance of these sets under the flow verifies stability, while their outward drift proves instability.
Level sets in classical dynamical systems are static geometric objects. The phase space itself, where level sets reside, was co-developed in the late nineteenth century by Boltzmann [3], Poincaré [13], and Gibbs [11]. Boltzmann’s 1872 paper introduced the distribution function on position-momentum space; Poincaré’s four-part memoir of 1881–1886 established the qualitative geometric theory of trajectories in the plane; Gibbs’s 1902 treatise systematized the ensemble formalism. In this framework, the level sets of any function associated with a dynamical system $\dot{x}=F(x)$ (conserved quantities, Lyapunov functions, or the vector field itself) are fixed structures. Trajectories move through them, but the level sets do not move.
The entire geometric theory of dynamical systems (invariant manifolds, stable and unstable manifolds, basins of attraction, separatrices) rests on this assumption, as codified in the foundational survey of Smale [15] and in the Russian school’s treatment of bifurcations [1, 2]. In the latter, bifurcations alter the topology of level sets at isolated parameter values, but always according to equations fixed in advance; between bifurcation points the level set structure is entirely permanent.
Remark 1.1.
This static property is not a technical limitation, but a foundational assumption. A classical dynamical system is a triple $(T,X,\Phi)$ where $\Phi:X\times T\to X$ is a fixed evolution rule. The level sets of any function associated with the system are therefore determined once and for all when $\Phi$ is specified.
2 Level Sets in the Active Element Machine
The Active Element Machine (AEM) [5], when executing a Universal Turing Machine (UTM) program $\eta$, physically realizes the boolean transition functions
$\eta_{k}:\{0,1\}^{3}\times\{0,1\}^{2}\;\longrightarrow\;\{0,1\},\quad k\in\{0,1,2,3,4,5\},$
as active element firing patterns. Each function $\eta_{k}$ has a fixed logical level set structure. For example, $\eta_{3}^{-1}\{1\}\;=\;$
$\bigl\{\,(111,00),\,(110,00),\,(110,01),\,(110,10),\,(101,00),\,(101,01),\,(101,10),$
$(100,00),\,(100,10),\,(011,01),\,(011,10),\,(010,11),\,(010,01),\,(010,00)\bigr\}$,
which is a property of the Minsky UTM program being executed.
What is not fixed is the physical realization of these level sets within the AEM. At each UTM computational step, quantum random bits $R_{0},R_{1},\ldots,R_{13}$ are generated by a quantum random number generator of the type described in [16]; see also [9] for a comprehensive review of such devices. Meta commands set_dynamic_C and set_dynamic_E then reconfigure the active elements $D_{0},\ldots,D_{13}$ and their connections (the physical objects that compute membership in $\eta_{3}^{-1}\{1\}$), based on which of the $R_{k}$ fired. The same logical level set is realized by a different spatio-temporal firing pattern at each computational step.
More precisely, for each UTM instruction $(q,\alpha)\in Q\times A$, an invertible boolean function is defined
$B_{(q,\alpha)}:\{0,1\}^{15}\;\longrightarrow\;\{0,1\}^{15}$
(shown in Table 4 of [6]) that maps the random firing pattern $(x_{0},\ldots,x_{13},b_{3})$ to the firing activity $(y_{0},\ldots,y_{13},p_{3})$ of the elements computing $\eta_{3}$. The choice of which $B_{(q,\alpha)}$ governs the realization at step $j$ depends on the current UTM instruction $I_{j}$ – that is, on the dynamical state of the computation itself.
3 A New Mathematical Object
The concept is stated more precisely.
Definition 3.1 (Dynamic Level Set).
Let $f:X\to\{0,1\}$ be a fixed boolean function with level sets $f^{-1}\{0\}$ and $f^{-1}\{1\}$. A dynamic level set decomposition of $f$ is a family of invertible boolean functions $\{B_{s}\}_{s\in S}$, indexed by a state space $S$, together with an incomputable sequence $\omega:\mathbb{N}\to\{0,1\}^{n}$ and a computable function $h:\mathbb{N}\to S$, such that at step $j$ the physical encoding of $f^{-1}\{1\}$ is determined by $B_{h(j)}(\omega(j))$ rather than by any fixed representation. The level set as a mathematical set is invariant; its physical instantiation is a dynamical variable.
Remark 3.1.
This captures what is constructed in Procedure 2  [6]. The logical level sets $\eta_{k}^{-1}\{1\}$ are invariant properties of the UTM program $\eta$. The physical realizations are dynamical variables of the AEM, evolving by a rule that depends on the current UTM instruction (computable via $h$) and on quantum random input (incomputable by the two quantum randomness axioms; see [16, 9]).
A dynamic level set is novel for the following reasons.
It is not a parametric family.
In a parametric family $F(x;\,\theta(t))$, the level sets $F^{-1}(c;\,\theta(t))$ sweep through a predetermined sequence as $\theta$ varies. Here, no predetermined sequence exists: the quantum randomness axioms guarantee that the sequence of level set realizations is itself Turing incomputable, so it cannot have been fixed in advance.
It is not bifurcation theory.
In bifurcation theory [1, 2], level set topology changes at isolated parameter values according to fixed equations of motion. In procedure 2  [6] , the physical realization changes at every computational step, not at exceptional points, and is governed by an incomputable process rather than a smooth family.
It is not a probabilistic Turing machine in the sense of de Leeuw et al.
The classical 1956 result shows that a Turing machine augmented with a random oracle with probability $p$, where $p$ is Turing computable, computes no more than a deterministic one. That result applies because the Turing machine program, and therefore its level set structure, does not change in response to random input. In the AEM, the Meta command reconfigures the physical realization of the level sets at each step based on the random input. The program’s logical content is invariant; its physical encoding is not. This is the structural distinction that allows this new model of computation to escape the 1956 result.
4 Incomputability
Lemma 4.1 of [6] is a fundamental lemma for proving incomputability.
Lemma 4.1 (Fiske 2012).
Let $\varphi:\mathbb{N}\to\{0,1\}$ be an incomputable function. For each $k$ with $1\leq k\leq m$, let $B_{k}:\{0,1\}^{n}\to\{0,1\}^{n}$ be an invertible boolean function, and let $h:\mathbb{N}\to\{1,\ldots,m\}$ be a computable function. Define $g:\mathbb{N}\to\{0,1\}$ by
$\bigl(g(jn+1),\,\ldots,\,g((j+1)n)\bigr)\;=\;B_{h(j)}\bigl(\varphi(jn+1),\,\ldots,\,\varphi((j+1)n)\bigr).$
Then $g$ is incomputable.
The proof is by contradiction: if $g$ were computable, then since each $B_{k}$ is invertible and $h$ is computable, one could recover $\varphi$ by applying $B_{h(j)}^{-1}$ to each $n$-tuple, contradicting the incomputability of $\varphi$.
The dynamical level set construction generates an incomputable firing pattern sequence whenever the source of randomness measures quantum events. The Turing incomputability of the observable behavior is a mathematical consequence of the new level set concept, not an additional assumption.
5 Connection to the Principle of Self-Modifiability
This concept links the 2012 Turing Centenary paper [6] and subsequent papers on self-modifiable dynamical systems [7, 8]. In the AEM executing the UTM:
  • The state space is fixed: the set of UTM configurations.
  • The logical transition function $\eta$ is fixed: it is a specific UTM program.
  • The physical realization of the level sets of $\eta$ – the actual dynamical objects computing membership in $\eta_{k}^{-1}\{0\}$ and $\eta_{k}^{-1}\{1\}$ – are self-modifying during execution.
This construction illustrates the Principle of Self-Modifiability [7] applied to the level set structure of the system. The system modifies the geometric objects that define how it computes, while it is computing. The level sets as logical objects are invariant; the level sets as physical dynamical objects are self-modifiable. This distinction between the logical and physical identity of a mathematical structure is itself a new conceptual contribution.
6 Level sets and Phase spaces
The phase space framework inherited from Boltzmann [3], Poincaré [13], and Gibbs [11] treats every geometric object in the phase portrait – including level sets of the Lyapunov function [12], and the stable and unstable manifolds identified in Smale [15] – as a permanent feature of the landscape. The bifurcation theory of the Russian school [1, 2] allows the topology of these objects to change, but only at isolated points and always according to equations fixed in advance. None of this allows the physical realization of level sets to be a dynamical variable of the system itself.
The closest standard concepts are mathematically distinct:
  • Non-autonomous systems: the vector field depends on time, but according to a rule fixed before evolution begins.
  • Variable structure control: the controller switches between pre-specified rule sets, but the full menu of rules is fixed before execution.
  • Randomized Turing machines: random input does not change program structure, as established in the 1956 result.
Dynamic level sets do not appear to have been formalized, or studied in the standard mathematical literature prior to the paper, Turing Incomputable Computation.
7 Summary
The construction in [6], implemented as Procedure 2, implicitly introduces a new mathematical object: a system in which the logical structure of level sets is invariant, but their physical realization is a dynamical variable, evolving under the self-modifiability of the meta command in the AEM and incomputable quantum randomness [16, 9]. This mathematical object:
  1. 1.
    Escapes the 1956 de Leeuw, et. al. results by using the mechanism of a Meta-command level set realization;
  2. 2.
    Generates Turing incomputable behavior from a finite program, as proved by Theorem 4.2 via Lemma 4.1 of [6];
  3. 3.
    Achieves Shannon perfect secrecy [14] of machine execution, as proved by Theorem 4.3 of [6]; and
  4. 4.
    Anticipates the Principle of Self-Modifiability developed in [7], providing its earliest and perhaps most mathematically explicit instance implemented with an active element machine [5].